OpenStudy (anonymous):

3/12 x -12 + (-8 - 3 ) - 12^2 explain/check answer.

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

-11 + -48-24?

OpenStudy (studygurl14):

\(\large\frac{3}{12}(-12)+(-8-3)-12^2\) Here's a clue...you can do this: \(\large\frac{3(-12)}{12}\rightarrow\frac{-3\cancel{12}}{\cancel{12}}\rightarrow=-3\)

OpenStudy (studygurl14):

And remember, \(\large 12^2 = (12)(12)\)

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

what about

OpenStudy (anonymous):

[(12 x 9)-1] ^2-3?

OpenStudy (studygurl14):

That would be \([(12\times9-1]^2-3\) Right?

OpenStudy (anonymous):

yeah

OpenStudy (studygurl14):

(Insert a parenthesis after the 9)

OpenStudy (studygurl14):

Okay, then...

OpenStudy (studygurl14):

\([(12\times9)-1]^2-3\rightarrow (108-1)^2-3\rightarrow (107)^2-3...\) Can you do the rest yourself?

OpenStudy (anonymous):

yeah let me try

OpenStudy (anonymous):

actually no i cant like my brain just froze when i saw the equation O.O

OpenStudy (studygurl14):

Do this: \(\large (107)(107)-3\)

OpenStudy (anonymous):

11446

OpenStudy (anonymous):

?